CO_{2}/MWSB GENERATION WITH HIGH-J, SEQUENCE, AND HOT BAND CO_{2} LASER LINES AND BROAD-BAND SCAN CAPABILITY

ZHEN-DONG SUN, QIANG LIU, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5; VLADIMIR DOROVSKIKH, M. YU. TRETYAKOV, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5, Permanent address: Institute of Applied Physics, Nizhny Novgorod, Russia 603950; R.M. LEES AND \underline{LI-HONG XU}, Department of Physical Sciences, University of New Brunswick, Saint John, NB, Canada E2L 4L5.

A broadband tunable CO_{2}-laser/microwave sideband spectrometer is now in operation at UNBSJ with the following noteworthy features: (i) The source coverage has been significantly improved as we have obtained useful sideband power for regular CO_{2} lines up to very high J (J_{max} = 56) as well as hot and sequence band CO_{2} laser lines, giving continuous coverage over significant spectral regions; (ii) The whole system has now been automated with sophisticated computer-controlled sweep and data capture in which the HP-MW synthesizer and Fabry-Perot etalon filter are scanned simultaneously to give wide-band continuous sweep over the full 7-18 GHz range for either sideband; (iii) Balanced detection of sample and reference has greatly improved the S/N ratio through compensation for background variations. Tests on the dense methanol spectrum in our room temperature multipass static cell around 10 μm have shown excellent resolution and S/N ratio using the hot and sequence band CO_{2} lines. Use of the source in the first tests of our supersonic slit-jet cooled molecular beam with OCS have shown significantly reduced line width compared to room-temperature static-cell values.